how many joules are required to heat 69g of water(liquid) from 69c to 96 c


1
Expert's answer
2022-03-22T23:29:03-0400

We apply the formula

q = mc"\Delta"T

Where m is the mass of water = 69/1000kg

c is the specific heat capacity of water = 4182J/kgoC

"\Delta"T is the change in temperature = 96-69=27oC


q = 0.069kg x 4182J/kgoC x 27oC = 7791.066J